OpenStudy (anonymous):

Use the technique of completing the square to transform the quadratic equation into the form (x + c)2 = a. 2x2 + 24x + 16 = 0

OpenStudy (anonymous):

start with \[x^2+12x+8=0\] because you need the leading coefficient to be 1

OpenStudy (anonymous):

then \[x^2+12x=-8\]and then \[(x+6)^2=-8+36\]

OpenStudy (anonymous):

(x + 6)2 = 28 <<---would that be the answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you.